DeGaetano Corporation manufactures a product which sells for $1,000. Standard variable production
costs are currently $450 per unit. Sales commissions total 20% of each unit sold. Fixed costs total
$495,000 for the year.

The company has projected that demand for their product will be limited to 2,000 units per year
for the in the upcoming year and foreseeable future. Due to extremely competitive market
forces, they cannot charge more than the current $1,000 price. The company wishes to
maintain a pre-tax profit of $315,000. Fixed costs are expected to increase by 10%. Per-unit
selling expenses are expected to remain unchanged. Using a target-costing approach, what
variable production cost per unit must the company attain in order to meet their profit objective?
(Round to the nearest penny)
Answer #1
Target sales revenue = 2000*1000 = $       20,00,000
Less: Sales commision at 20% $         4,00,000
Net sales revenue $       16,00,000
Less: Target pre tax profit $         3,15,000
Less: Expected fixed costs = 495000*110% = $         5,44,500
Target total variable cost $         7,40,500
Number of units 2000
Target variable cost per unit $             370.25